OpenStudy (masterchief):

The product of two numbers whose difference is 25. I'm not looking for a pair of numbers, I need the equation, apparently x^2-25 was wrong.

OpenStudy (anonymous):

xy x-y=25 --> y=x-25 so the product is P = xy = x(x-25) = x^2 - 25x

OpenStudy (masterchief):

Yes that was one of the choices on the handout, but I don't understand why the constant needed a variable too?

OpenStudy (mertsj):

If the difference of two numbers is 25, we can write x-y=25. That means that y is x-25 So the two numbers are x and x-25 The product of x and x-25 is x(x-25) Using the distributive property we have x^2-25x

OpenStudy (masterchief):

Ok then, thank you both

OpenStudy (mertsj):

yw